OpenStudy (anonymous):

if n/5-3n/10=1/5,what does n=?

OpenStudy (anonymous):

show work please so I can understand this problem better

OpenStudy (anonymous):

n=-2

OpenStudy (anonymous):

\[(n \div5) - (3n \div10)=(1\div5)\] \[(2n \div10)-(3n \div10) = (2\div10)\] \[2n -3n= 2\] \[-n=2\]

OpenStudy (anonymous):

so \[n=-2\]